A man standing at the top of a building throws a ball vertically upward with a velocity of 10 ms– 1 . The ball reaches the ground 4 s later. What is the maximum height reached by the ball? How high is the building
 

Grade:11

1 Answers

Ananya
2 years ago
Since time taken to reach ground=4s
Time taken to reach max hgt.=2s
We know s=ut+1/2at²
H(max)=10*2-1/2*9.8*2*2
              =20-19.6=0.4m